### v1

@@ 4 Numerical experiments
In this section, we test the infinite-dimensional Nyström approximation proposed in this work to compute low-rank approximations to different integral kernel operators. Algorithm 1 presents the pseudocode of a variant of [28, Alg. 16] for non-negative self-adjoint trace-class operators. We emphasize that the orthonormalization steps in line 2 and the shift in line 5 improve the numerical stability of the algorithm; see e.g. [39]. Moreover, our implementation of the Nyström approximation computes the Cholesky factorization of the symmetric part of $\bm{\Omega}^{*}\bm{Y}_{\nu}$ in line 6. Finally, we note that in exact arithmetic, with $\nu=0$ in line 4, the approximation returned by Algorithm 1 is mathematically equivalent to Eq. 20.
The aim of our numerical experiments is to validate Algorithm 1. For this purpose, we consider an interval $D=[a,b]$ and carry out all operations on functions on $D$ (approximately) using the Chebfun software package [15], which is an open-source MATLAB package containing algorithms for performing numerical linear algebra operations on functions, such as the QR decomposition [36, 37]. In all experiments, we select the squared-exponential kernel:
$K(x,y)=\exp\left(-\frac{2(x-y)^{2}}{(b-a)^{2}\ell^{2}}\right),\quad x,y\in[a,b],$
as covariance kernel associated with the random field $\Omega$ in Algorithm 1. We vary the length-scale parameter $\ell$ across three values: $\ell=1,0.1,0.01$. A smaller value for $\ell$ results in smoother Gaussian processes that are more biased towards certain spatial directions. Conversely, a larger value of $\ell$ results in rougher Gaussian processes that are less spatially biased. We vary $\ell$ to investigate the effect of the smoothness of the Gaussian process on the result of the low-rank approximation.
@@ 4.1 A pretty function
In this first example, we compute an approximation to the integral operator defined by the kernel [35]
$G(x,y)=\frac{1}{1+100(x^{2}-y^{2})^{2}},\quad x,y\in[-1,1],$ (27)
@@
Figure 1: (a) Exact kernel defined by Eq. 27 along with convergence of the Nyström approximation for different values of $\ell$ (b). (c)-(e) Rank-$40$ Nyström approximations of the kernel for $\ell=1,0.1,0.01$, respectively.
@@ 4.2 Matérn Kernels
In this second example, we approximate the integral operator defined by the Matérn-$1/2,3/2,$ and $5/2$ kernels [34, Chapt. 4]
$G_{1/2}(x,y)=\exp(-|x-y|),\quad x,y\in[-1,1];$ (28)
$G_{3/2}(x,y)=(1+\sqrt{3}|x-y|)\exp\left(-\sqrt{3}|x-y|\right),\quad x,y\in[-1,1];$ (29)
$G_{5/2}(x,y)=\left(1+\sqrt{5}|x-y|+\frac{5}{3}(x-y)^{2}\right)\exp\left(-\sqrt{5}|x-y|\right),\quad x,y\in[-1,1].$ (30)
The Matérn class is an important class of covariance kernels, frequently appearing in machine learning. The parameter $\nu=1/2,3/2,5/2$ determines the spectral decay of the kernel and thus the smoothness of the Gaussian process, with higher $\nu$ implying faster decay and smoother Gaussian processes. The results are presented in Fig. 2 for $G_{1/2},G_{3/2},$ and $G_{5/2}$, respectively.
@@
Figure 2: The contour plots show the exact kernel defined by (28) and the approximated kernels using rank $100$ Nyström approximations. The error plots show the relative error in the Hilbert-Schmidt norm. Optimal denotes the best low-rank approximation error.
4.3 Green’s function for an elliptic differential operator
In this example we consider the operator $\mathcal{A}_{\lambda}$ so that $u=\mathcal{A}_{\lambda}f$ solves the equation
$-\Delta u(\bm{x})+\lambda u(\bm{x})=f(\bm{x}),\quad\bm{x}\in[0,2\pi]^{d},$
where $\lambda\geq 0$, with Dirichlet boundary conditions and $d\in\{1,2,3\}$ the spatial dimension. Let $\mathcal{A}_{0}$ have Hilbert-Schmidt decomposition
$[\mathcal{A}_{0}f](\bm{x})=\sum\limits_{i=1}^{\infty}\sigma_{i}u_{i}(\bm{x})\langle u_{i},f\rangle,$
@@ 4.3 Green’s function for an elliptic differential operator
$[\mathcal{A}_{\lambda}f](\bm{x})=\sum\limits_{i=1}^{\infty}g_{\lambda}(\sigma_{i})u_{i}(\bm{x})\langle u_{i},f\rangle=[g_{\lambda}(\mathcal{A}_{0})f](\bm{x}),\quad g_{\lambda}(\sigma)=\frac{\sigma}{\lambda\sigma+1}.$
For each $\lambda\geq 0$ the function $g_{\lambda}$ is operator monotone [5, Sec. V]. Recent work has shown that if $\widehat{\mathcal{A}}$ is a near-optimal low-rank approximation to $\mathcal{A}$, then $g(\widehat{\mathcal{A}}_{0})$ is a near-optimal low-rank approximation to $g(\mathcal{A})$, for any operator monotone function $g$ satisfying $g(0)=0$; see [32, 33]. Hence, to obtain a low-rank approximation to $\mathcal{A}_{\lambda}$ we compute a Nyström approximation $\widehat{\mathcal{A}}_{0}$ to $\mathcal{A}_{0}$ and then approximate $\mathcal{A}_{\lambda}$ with $g_{\lambda}(\widehat{\mathcal{A}}_{0})$. In our experiments, we set $\lambda=1$ and $d=1$. In this case, the Green’s function is given by
$G(x,y)=\min(x,y)-\frac{xy}{2\pi}.$ (31)
The results are displayed in Fig. 3.
@@
Figure 3: The contour plots show the exact kernel defined by (31) and the approximated kernels using rank $100$ Nyström approximations. The error plots show the relative error in the Hilbert-Schmidt norm. Optimal denotes the best low-rank approximation error.
@@ 5 Conclusions
In this work, we have presented an infinite-dimensional extension of the randomized Nyström approximation. We first established bounds for the finite-dimensional Nyström approximation when the columns of the sketch matrix $\bm{\Omega}$ are drawn independently from a non-standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{K})$. Subsequently, through a continuity argument, we provided analogous bounds for the infinite-dimensional extension of the Nyström approximation for Hilbert-Schmidt operators. Additionally, in the process of analyzing the Nyström approximation for Hilbert-Schmidt operators, we have also improved the existing bounds for the randomized SVD for Hilbert-Schmidt operators.
@@ Appendix A Properties of Gaussian matrices
Lemma A.1.
Let $\bm{C}\in\mathbb{R}^{m_{1}\times m_{2}}$, $\bm{D}\in\mathbb{R}^{n_{1}\times n_{2}}$ be two matrices, consider an $m_{2}\times n_{1}$ standard Gaussian matrix $\bm{\Psi}$, and define $\bm{\Phi}=\bm{C}\bm{\bm{\Psi}}\bm{D}$. Then, $\mathbb{E}_{\bm{\Psi}}[\bm{\Phi}\bm{\Phi}^{*}\bm{\Phi}]=\bm{0}$.
Proof A.2.
First note that the expectation exists, since each entry of $\bm{\Phi}\bm{\Phi}^{*}\bm{\Phi}$ is a linear combination of products of Gaussian random variables, which always has a finite expectation. The distribution of $\bm{\Psi}$ and, in turn, $\bm{\Phi}$ is symmetric, which gives
$-\mathbb{E}_{\bm{\Psi}}[\bm{\Phi}\bm{\Phi}^{*}\bm{\Phi}]=\mathbb{E}_{\bm{\Psi}}[-\bm{\Phi}\bm{\Phi}^{*}\bm{\Phi}]=\mathbb{E}_{\bm{\Psi}}[(-\bm{\Phi})(-\bm{\Phi})^{*}(-\bm{\Phi})]=\mathbb{E}_{\bm{\Psi}}[\bm{\Phi}\bm{\Phi}^{*}\bm{\Phi}],$
implying the result of the lemma.
@@
$\mathbb{E}\big{[}\|\bm{B}+\bm{C}\bm{\bm{\Psi}}\bm{D}\|_{\textup{F}}^{2}\big{]}$ $=\|\bm{B}\|_{\textup{F}}^{2}+\|\bm{C}\|_{\textup{F}}^{2}\|\bm{D}\|_{\textup{F}}^{2},$ (32a)
$\mathbb{E}\big{[}\|\bm{C}\bm{\Psi}\bm{D}\|_{(4)}^{4}\big{]}$ $=\|\bm{C}\|_{(4)}^{4}\|\bm{D}\|_{(4)}^{4}+\|\bm{C}\|_{\textup{F}}^{4}\|\bm{D}\|_{(4)}^{4}+\|\bm{C}\|_{(4)}^{4}\|\bm{D}\|_{\textup{F}}^{4},$ (32b)
$\mathbb{E}\big{[}\|\bm{C}\bm{\bm{\Psi}}\bm{D}\|_{2}^{2}\big{]}$ $\leq\left(\|\bm{C}\|_{\textup{F}}\|\bm{D}\|_{2}+\|\bm{C}\|_{2}\|\bm{D}\|_{\textup{F}}\right)^{2},$ (32c)
$\mathbb{E}\big{[}\|\bm{C}\bm{\Psi}\bm{D}\|_{\textup{F}}^{4}]$ $=2\|\bm{C}\|_{(4)}^{4}\|\bm{D}\|_{(4)}^{4}+\|\bm{C}\|_{\textup{F}}^{4}\|\bm{D}\|_{\textup{F}}^{4}.$ (32d)
Proof A.4.
@@
[…]
@@
Let $\bm{\Omega}_{1}\in\mathbb{R}^{k\times(k+p)}$ be a random matrix whose columns are i.i.d. $\mathcal{N}(0,\widetilde{\bm{K}}_{11})$ and let $\bm{B}\in\mathbb{R}^{k\times n}$ be an arbitrary matrix. Then, the following relation hold for $k\geq 1$ and $p\geq 2$:
$\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\bm{B}\|_{\textup{F}}^{2}]$ $=\frac{\operatorname{tr}(\bm{B}^{*}\widetilde{\bm{K}}_{11}^{-1}\bm{B})}{p-1}=\frac{\|\bm{B}^{*}\widetilde{\bm{K}}_{11}^{-1/2}\|_{\textup{F}}^{2}}{p-1},$ (33a)
@@
$\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}]$ $\leq\frac{e^{2}(k+p)}{(p-1)(p+1)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}.$ (34a)
@@
$\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}]$ $=\frac{(p-1)\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}+\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})^{2}}{p(p-1)(p-3)}\leq k\frac{k+p-1}{p(p-1)(p-3)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2},$ (35a)
$\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}]$ $=\frac{(p-2)\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})^{2}+2\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}}{p(p-1)(p-3)}\leq k\frac{kp-2k+2}{p(p-1)(p-3)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2}.$ (35b)
Proof A.6.
First, Eq. 33a is proven in [13, Lem. 3.12] and Eq. 34a follows from in [30, Lem. 3.1] and the fact $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}]\leq\mathbb{E}[\|\bm{\Psi}^{\dagger}\|_{2}^{2}]\|\bm{K}_{11}^{-1}\|_{2}$ where $\bm{\Psi}$ is a $k\times(k+p)$ standard Gaussian matrix. We proceed to prove the equality in Eq. 35a and introduce the random variable $\bm{X}=\bm{\Omega}_{1}\bm{\Omega}_{1}^{*}\sim\mathcal{W}_{k}(\widetilde{\bm{K}}_{11},k+p)$, such that
@@
where the last equality follows from [23, Thm. 2.4.14]. The equality in Eq. 35b follows from the fact that $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}]=\mathbb{E}[\operatorname{tr}(\bm{X}^{-1})^{2}]=\mathbb{E}[\operatorname{tr}\left(\bm{X}^{-1}\otimes\bm{X}^{-1}\right)]$. We then exploit a relation on $\mathbb{E}[\bm{X}^{-1}\otimes\bm{X}^{-1}]$ [23, Thm. 2.4.14] to obtain
@@
The next lemma is a generalization and a consequence of [20, Prop. 10.4], which provides tailbounds on the Frobenius and spectral norms of pseudoinverted standard Gaussian matrices.

### v2

@@ 4 Numerical experiments
In this section, we test the infinite-dimensional Nyström approximation proposed in this work to compute low-rank approximations to different trace class operators. Algorithm 1 presents the pseudocode of a variant of [30, Alg. 16] for non-negative self-adjoint trace-class operators. We emphasize that the orthonormalization steps in line 2 and the shift in line 5 improve the numerical stability of the algorithm; see e.g. [40]. Moreover, our implementation of the Nyström approximation computes the Cholesky factorization of the symmetric part of $\bm{\Omega}^{*}\bm{Y}_{\nu}$ in line 6. Finally, we note that in the ideal setting when we are able to compute with functions exactly, with $\nu=0$ in line 4, the approximation returned by Algorithm 1 is mathematically equivalent to Eq. 23.
The aim of our numerical experiments is to validate Algorithm 1. For this purpose, we consider a set $D=[-1,1]^{d}$ for $d=1,2$ and compute low-rank approximations to different operators $\mathcal{A}:L^{2}(D)\to L^{2}(D)$. In most of the experiments, we will carry out all operations on functions on $D$ (approximately) using the Chebfun software package [16], which is an open-source MATLAB package containing algorithms for performing numerical linear algebra operations on functions and quasimatrices, such as the QR decomposition [37, 38]. A common choice for the covariance kernel associated with the random field $\Omega$ in Algorithm 1 is the squared-exponential kernel:
(30) $K(\bm{x},\bm{y})=\prod\limits_{i=1}^{d}\exp\left(-\frac{(\bm{x}_{i}-\bm{y}_{i})^{2}}{2\ell^{2}}\right),\quad\bm{x},\bm{y}\in[-1,1]^{d}.$
A large value for the length-scale parameter $\ell$ results in smoother Gaussian processes that are more biased towards certain spatial directions. Conversely, a smaller value of $\ell$ results in rougher Gaussian processes that are less spatially biased. We will vary $\ell$ to investigate the effect of the smoothness of the Gaussian process on the result of the low-rank approximation.
@@ 4.1 A pretty function
In this first example, we compute an approximation to the integral operator $\mathcal{A}f=\int_{-1}^{1}G(\cdot,y)f(y)\,\mathrm{d}y$ defined by the kernel [36]
(31) $G(x,y)=\frac{1}{1+100(x^{2}-y^{2})^{2}},\quad x,y\in[-1,1],$
@@
[…]
@@ 4.2 Matérn Kernels
In this second example, we consider the integral operator $\mathcal{A}f=\int_{-1}^{1}G(\cdot,y)f(y)\,\mathrm{d}y$ defined by the Matérn-$1/2$, $3/2$, and $5/2$ kernels [35, Chapt. 4]:
(32) $G_{1/2}(x,y)=\exp(-|x-y|),\quad x,y\in[-1,1];$
(33) $G_{3/2}(x,y)=(1+\sqrt{3}|x-y|)\exp\left(-\sqrt{3}|x-y|\right),\quad x,y\in[-1,1];$
(34) $G_{5/2}(x,y)=\left(1+\sqrt{5}|x-y|+\frac{5}{3}(x-y)^{2}\right)\exp\left(-\sqrt{5}|x-y|\right),\quad x,y\in[-1,1].$
The Matérn class is an important class of low-regularity covariance kernels, frequently appearing in machine learning, where a common task is to compute low-rank approximations of discretizations of the associated kernels. A key benefit of the proposed framework in this work is that once a low-rank approximation to the continuous kernel function is computed, one can obtain a low-rank approximation to any discretization of the kernel through a discretization of the low-rank approximation.
In this example, we are going to demonstrate how one can incorporate prior knowledge of the kernel into the random fields to obtain more accurate low-rank approximations.
@@
Figure 2: The contour plots (a),(c), and (e) show the Matérn kernels defined in (32)-(34). The error plots (b) and (d) show the relative error in the trace norm when approximating $G_{1/2}$ and $G_{5/2}$, respectively, when using the squared exponential kernel in (30) with $\ell=1,0.01$ and the Matérn-3/2 kernel in (33) as covariance kernels for the random fields. Optimal denotes the best low-rank approximation error. The plots in (f) show sample paths of the different Gaussian processes. The paths have been normalized so that the maximum absolute value is equal to 1.
@@
Figure 3: Figure comparing the numerical results against the theoretical upper bound in (26b) when approximating the Matérn-5/2 kernel using the squared-exponential with length parameters $\ell=1$ and $\ell=0.01$ and the Matérn-3/2 kernel as covariance kernels. Optimal denotes the best low-rank approximation error. The full line shows the numerical results. The dashed lines shows the theoretical upper bound in (26b).
We compute a Nyström approximation for $G_{1/2}$ and $G_{5/2}$ using the squared-exponential kernel with length-scale parameters $\ell=1$ and $\ell=0.01$, as well as $G_{3/2}$ as a covariance kernel for the random fields. Using $G_{3/2}$ as a covariance kernel incorporates prior knowledge of the operator’s eigenfunctions, since the Matérn-$\nu$ kernels vary continuously with $\nu$. The results are presented in Figure 2, which demonstrate that incorporating such prior knowledge can lead to more accurate low-rank approximations. For the Matérn-5/2 kernel, we have also included a figure comparing the numerical results with the upper bound in (26b); see Figure 3. Notably, from Figure 2 and Figure 3, we observe that using $G_{3/2}$ as a covariance kernel for random fields yields excellent results. For the squared-exponential kernel, when $\ell$ is small, there is significantly less variance in the directions of $U_{2}$, i.e., $\beta_{k}^{(\mathrm{Tr})}$ is small for the corresponding random field. This results in near-optimal low-rank approximations for small values of $k$. However, as $k$ increases, the random fields become too spatially biased, leading to large values of $\delta_{k}^{(k)}$. In contrast, when $\ell$ is small, the random fields exhibit a higher variance in all directions due to less spatial bias. This leads to larger values of $\beta_{k}^{(\mathrm{Tr})}$ for small $k$, but as higher-frequency terms are needed to capture the eigenfunctions associated with smaller eigenvalues, the approximation becomes more efficient for larger $k$.
In cases where no prior knowledge of the operator is available, we recommend using the roughest possible random fields, subject to computational constraints. Although sampling from rougher random fields is typically more computationally expensive than sampling from smoother ones, since the corresponding Karhunen-Loève expansion converges slower, they offer a safer choice, as they are less spatially biased.
4.3 Approximating a Karhunen-Loève expansion
In our next example we consider the task of sampling from a Gaussian process. Given a symmetric positive semi-definite kernel $G:D\times D\to\mathbb{R}$, inducing an integral operator $\mathcal{A}f=\int_{D}G(\cdot,y)f(y)\,\mathrm{d}y$, a common task in computational statistics is to sample from $\mathcal{GP}(0,G)$. This can be achieved in a few ways. For example, if the Mercer expansion of $G$ is known, one can sample from $\mathcal{GP}(0,G)$ through the Karhunen-Loève expansion (22):
$\omega=\sum\limits_{i=1}^{\infty}\lambda_{i}^{1/2}\zeta_{i}u_{i}\sim\mathcal{GP}(0,G),$
where $\zeta_{1},\zeta_{2},\cdots\sim\mathcal{N}(0,1)$ are mutually independent. However, computing a Mercer expansion of $G$ can be computationally intractable. One can also sample from a discretized Gaussian process by sampling from $\mathcal{N}(\bm{0},\bm{G})$, where $\bm{G}_{ij}=G(x_{i},x_{j})$ and $\{x_{i}\}_{i=1}^{n}$ are the discretization points. This requires computing a square-root or a Cholesky factorization of $\bm{G}$, which costs $O(n^{3})$ operations and becomes prohibitively expensive for large $n$.
If we have access to an accurate Nyström approximation $G_{k}$ to $G$ one can easily sample from $\mathcal{GP}(0,G_{k})$ as
$\widehat{\omega}=\sum\limits_{i=1}^{k}\widehat{\lambda}_{i}^{1/2}\widehat{\zeta}_{i}\widehat{u}_{i}\sim\mathcal{GP}(0,G_{k}),$
where $(\widehat{\lambda}_{i},\widehat{u}_{i})$ are the eigenpairs of $G_{k}$ for $i=1,\ldots,k$, and $\widehat{\zeta}_{1},\ldots,\widehat{\zeta}_{k}\sim\mathcal{N}(0,1)$ are independent. Computing a Nyström approximation of $\mathcal{G}$ is usually significantly less expensive than computing a Mercer expansion of $G$. Furthermore, choosing $\widehat{\zeta}_{i}$ so that $\mathbb{E}[\zeta_{i}\widehat{\zeta}_{j}]=\langle u_{i},\widehat{u}_{j}\rangle_{L^{2}(D)}$ yields $\mathbb{E}\|\omega-\widehat{\omega}\|_{L^{2}(D)}^{2}\leq\|G-G_{k}\|_{\mathrm{Tr}}$. [^3] This fact requires that $G-G_{k}$ is a positive semi-definite kernel, which is true whenever $G_{k}$ is the kernel of the Nyström approximation. However, this fact is not true for arbitrary approximations. Hence, $\mathcal{W}_{2}^{2}(\mathcal{GP}(0,G),\mathcal{GP}(0,G_{k}))\leq\|G-G_{k}\|_{\mathrm{Tr}}$, where $\mathcal{W}_{2}$ denotes the Wasserstein distance. Therefore, as long as $G_{k}$ is an accurate low-rank approximation to $G$, $\mathcal{GP}(0,G_{k})$ is a good approximation to $\mathcal{GP}(0,G)$ in the Wasserstein distance.
@@ 4.3 Approximating a Karhunen-Loève expansion
Figure 4: The left contour plot shows an exact sample from $\mathcal{GP}(0,G)$ and the left contour plot shows a sample from $\mathcal{GP}(0,G_{30})$ so that the mean-squared error equals $\|G-G_{30}\|_{\mathrm{Tr}}$. The error plots show the relative error in the trace norm. Optimal denotes the best low-rank approximation error.
In this example we consider $G$ to be the squared-exponential kernel (30) with $d=2$ and $\ell=0.4$. The random fields are drawn from $\mathcal{GP}(0,K)$ with $K(\bm{x},\bm{y})=\sum_{i=1}^{625}p_{i}(\bm{x})p_{i}(\bm{x})$ where $\{p_{1},\ldots,p_{625}\}$ is an orthonormal basis obtained by taking the tensor product of the first 25 Legendre polynomials. For the purpose of this experiment, we discretize the squared-exponential kernel by expanding it onto the orthonormal basis obtained by taking the tensor product of the first $100$ Legendre polynomials. The results are displayed in Fig. 4.
4.4 An example from Bayesian inverse problems
Our final example comes from A-optimal design of Bayesian inverse problems. In this setting, we have a linear parameter-to-observable map $\mathcal{F}:L^{2}(D)\to\mathbb{R}^{n}$, which maps, for example, the initial condition of a partial differential equation to the function values at a set of discrete points of the solution. In the context of A-optimal design of infinite-dimensional linear inverse problems one wants to compute a low-rank approximation of the operator
$\mathcal{A}=\mathcal{C}_{0}^{1/2}\mathcal{F}^{*}\bm{\Sigma}_{\text{noise}}\mathcal{F}\mathcal{C}_{0}^{1/2},$
where $\mathcal{C}_{0}$ is the covariance operator of the Gaussian prior distribution and $\bm{\Sigma}_{\text{noise}}$ is the covariance matrix of measurement errors [2, 23].
As performed in [2, 23], we let $\mathcal{C}_{0}$ be the solution operator of the Biharmonic equation on $\Omega=[-1,1]^{2}$ with Dirichlet boundary conditions, $\bm{\Sigma}_{\text{noise}}\propto\bm{I}$, and denote by $\mathcal{F}$ the operator that maps the (space-dependent) initial condition $\theta\in L^{2}(\Omega)$ to the time-dependent solution $u$ of the following parabolic differential equation:
(35) $\begin{split}\partial_{t}u=\Delta u+0.1u\quad\text{in }\Omega\times\mathbb{R}_{\geq 0},\\ u=0\quad\text{on }\partial\Omega,\quad\text{and}\quad u(\cdot,0)=\theta\quad\text{in }\Omega,\end{split}$
on a $50\times 50$ equispaced grid of $\Omega$ and at times $t=0.1,0.5,1$.
For the purpose of this experiment we solve (35) by first discretizing in space using finite differences on a $50\times 50$ equispaced grid, yielding an ordinary differential equation (ODE). We then proceed in time by solving the resulting ODE using exponential integrators. We let the covariance kernel for the random fields be the solution operator of the Biharmonic equation on $[-1,1]^{2}$ with Direchlet boundary conditions. The results are displayed in Fig. 5, which demonstrates that the Nyström approximation yields near-optimal low-rank approximations to the operator.
[…]
Figure 5: The contour plots shows the solution of (35) for different times. The error plots show the relative error in the trace norm norm. Optimal denotes the best low-rank approximation error.
@@ 5 Conclusions
In this work, we have presented an infinite-dimensional extension of the randomized Nyström approximation. We first established bounds for the finite-dimensional Nyström approximation when the columns of the sketch matrix $\bm{\Omega}$ are drawn independently from a non-standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{K})$. Subsequently, through a continuity argument, we provided analogous bounds for the infinite-dimensional extension of the Nyström approximation for Hilbert-Schmidt operators. Additionally, in the process of analyzing the Nyström approximation for Hilbert-Schmidt operators, we have also improved the existing bounds for the randomized SVD for Hilbert-Schmidt operators. Our analysis is complemented by a number of numerical experiments on the application of the proposed infinite-dimensional Nyström method and demonstrate its potential in Gaussian process sampling or Bayesian inverse problems.
@@ Appendix A Properties of Gaussian matrices
(36a) $\mathbb{E}\big{[}\|\bm{B}+\bm{C}\bm{\bm{\Psi}}\bm{D}\|_{\textup{F}}^{2}\big{]}$ $=\|\bm{B}\|_{\textup{F}}^{2}+\|\bm{C}\|_{\textup{F}}^{2}\|\bm{D}\|_{\textup{F}}^{2},$
(36b) $\mathbb{E}\big{[}\|\bm{C}\bm{\Psi}\bm{D}\|_{(4)}^{4}\big{]}$ $=\|\bm{C}\|_{(4)}^{4}\|\bm{D}\|_{(4)}^{4}+\|\bm{C}\|_{\textup{F}}^{4}\|\bm{D}\|_{(4)}^{4}+\|\bm{C}\|_{(4)}^{4}\|\bm{D}\|_{\textup{F}}^{4},$
(36c) $\mathbb{E}\big{[}\|\bm{C}\bm{\bm{\Psi}}\bm{D}\|_{2}^{2}\big{]}$ $\leq\left(\|\bm{C}\|_{\textup{F}}\|\bm{D}\|_{2}+\|\bm{C}\|_{2}\|\bm{D}\|_{\textup{F}}\right)^{2},$
(36d) $\mathbb{E}\big{[}\|\bm{C}\bm{\Psi}\bm{D}\|_{\textup{F}}^{4}]$ $=2\|\bm{C}\|_{(4)}^{4}\|\bm{D}\|_{(4)}^{4}+\|\bm{C}\|_{\textup{F}}^{4}\|\bm{D}\|_{\textup{F}}^{4}.$
Proof A.2.
@@
where we combined [21, Prop. 10.1] with the equality $\mathbb{E}[\operatorname{tr}(\bm{B}^{*}\bm{\Phi})]=0$ since $\bm{\Psi}$ has zero mean. A similar result is found in [14, Lem. 3.11]. The equality Eq. 36b and inequality Eq. 36c can be found in [39, Lem. B.1] and [18, Prop. B.3], respectively. We now conclude with the proof of (36d). Let $\bm{E}=\bm{D}^{*}\otimes\bm{C}$ and $\bm{\psi}=\text{vec}(\bm{\Psi})$. Then,
@@
Let $\bm{\Omega}_{1}\in\mathbb{R}^{k\times(k+p)}$ be a random matrix whose columns are i.i.d. $\mathcal{N}(0,\widetilde{\bm{K}}_{11})$ and let $\bm{B}\in\mathbb{R}^{k\times n}$ be an arbitrary matrix. Then, the following relation holds for $k\geq 1$ and $p\geq 2$:
(37a) $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\bm{B}\|_{\textup{F}}^{2}]$ $=\frac{\operatorname{tr}(\bm{B}^{*}\widetilde{\bm{K}}_{11}^{-1}\bm{B})}{p-1}=\frac{\|\bm{B}^{*}\widetilde{\bm{K}}_{11}^{-1/2}\|_{\textup{F}}^{2}}{p-1},$
@@
(38a) $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}]$ $\leq\frac{e^{2}(k+p)}{(p-1)(p+1)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}.$
@@
(39a) $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}^{4}]$ $=\frac{(p-1)\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}+\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})^{2}}{p(p-1)(p-3)}\leq k\frac{k+p-1}{p(p-1)(p-3)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2},$
(39b) $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}]$ $=\frac{(p-2)\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})^{2}+2\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}^{2}}{p(p-1)(p-3)}\leq k\frac{kp-2k+2}{p(p-1)(p-3)}\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}^{2}.$
Proof A.4.
First, Eq. 37a is proven in [14, Lem. 3.12] and Eq. 38a follows from in [32, Lem. 3.1] and the fact $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{2}^{2}]\leq\mathbb{E}[\|\bm{\Psi}^{\dagger}\|_{2}^{2}]\|\bm{K}_{11}^{-1}\|_{2}$ where $\bm{\Psi}$ is a $k\times(k+p)$ standard Gaussian matrix. We proceed to prove the equality in Eq. 39a and introduce the random variable $\bm{X}=\bm{\Omega}_{1}\bm{\Omega}_{1}^{*}\sim\mathcal{W}_{k}(\widetilde{\bm{K}}_{11},k+p)$, such that
@@
where the last equality follows from [25, Thm. 2.4.14]. The equality in Eq. 39b follows from the fact that $\mathbb{E}[\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}^{4}]=\mathbb{E}[\operatorname{tr}(\bm{X}^{-1})^{2}]=\mathbb{E}[\operatorname{tr}\left(\bm{X}^{-1}\otimes\bm{X}^{-1}\right)]$. We then exploit a relation on $\mathbb{E}[\bm{X}^{-1}\otimes\bm{X}^{-1}]$ [25, Thm. 2.4.14] to obtain
@@
The next lemma is a generalization and a consequence of [21, Prop. 10.4], which provides tail bounds on the Frobenius and spectral norms of pseudoinverted standard Gaussian matrices.